# Where Do People Tell Stories Online? Story Detection Across Online Communities

### v2

@@ 6.1 What are the signs of storytelling spans?
| Measure | Effect Size () | Direction |
-value |
expert-annotated events |
2.122*** |
story | |
past tense |
1.742*** |
story | |
realis events |
1.655*** |
story | |
1st-person singular pronouns |
1.051*** |
story | |
3rd-person singular pronouns |
0.459*** |
story | |
entity mentions |
0.346*** |
story | |
concreteness |
0.329*** |
story |
0.001 |
non-past tense |
1.296*** |
non-story | |
is comment (vs. post) |
0.612*** |
non-story | |
2nd-person pronouns |
0.551*** |
non-story | |
sentence length |
0.345 |
| 1st-person plural pronouns |
0.345 |
Based on prior literature and our iterations of annotation and codebook refinement, we identify a set of features that we expect may be associated with storytelling in social media. These include entity and pronoun rates (Eisenberg and Finlayson, 2017; Piper and Bagga, 2022), events Hühn (2009); Gius and Vauth (2022); Sap et al. (2022); Sims et al. (2019), verb tense, concreteness (Piper and Bagga, 2022; Brysbaert et al., 2013), whether a text is a post or comment, and average sentence length. See Appendix A.9 for context on prior work related to these features, the precise definitions we adopted for this study, and a feature comparison test with another narrative detection dataset composed of mostly literary texts.
@@ 6.2 Where do people tell stories?
When interpreting these results, it is important to keep in mind that our comparisons are only across texts in the Webis-TLDR-17 dataset, i.e. coherent texts with summaries. Since many subreddits predominately include photos or other content, our storytelling rates should be interpreted as relative rankings rather than as absolute rates of storytelling. We provide additional results in Figure 5 in the Appendix that are not restricted to texts with summaries, validating that the overall pattern of our rankings holds when sampling over the full space of subreddit posts.
@@ 6.3 How distinctive are stories by community?
In Figure 2, we map communities across two axes: their story rate, predicted by our story detector, and the distinctiveness of their vocabulary. Categories of subreddits form interpretable clusters. We show 291 subreddits (those matching our filtering criteria, see §5), and Table 9 in the Appendix shows the “corners” of this plot, i.e., the subreddits with the most and least storytelling and distinctiveness. For example, subreddits in the stories category, such as r/Glitch_in_the_Matrix, tend to have both high rates of storytelling and low distinctiveness — these subreddits elicit stories that do not use consistently distinctive language. Some categories contain a wide range, e.g., the technology category contains subreddits varying across half the distinctiveness range, from r/apple (not at all distinctive) to r/buildapc (very distinctive stories). In Figure 6 in the Appendix, we show the full ranking of subreddit categories, which can aid researchers in determining whether to build a custom detector for their target domain or rely on a mixture of finetuning data to align with the diverse storytelling in some categories.
@@ 7 Case Study: r/ChangeMyView
Figure 3 shows the post topics that elicit comments with more or less storytelling and whose storytelling comments are more or less likely to receive a winning “delta” point, indicating that the poster was convinced by the comment’s argument. Topics are trained on the posts via a latent Dirichlet allocation (LDA) (Blei et al., 2003) model (), and after examination of the highest probability words and documents, we assign each topic a descriptive name. The y-axis shows the difference between the mean post topic probabilities for storytelling and non-storytelling comments, while the x-axis shows the difference between the proportion of winning comments that include storytelling and the proportion of non-winning comments that include storytelling. More details about our methods and the topic model are given Appendix A.3.
@@ 8 Discussion
The features that are most strongly associated with storytelling support prior work’s emphasis on storytelling’s basis in agent-centered, event-driven forms of communication that grounded in concrete settings (known as the “deictic theory” of storytelling; Piper and Bagga, 2022). However, unlike prior work, we find significantly more usage of first-person singular pronouns as well as significantly more storytelling in posts in contrast to comments.
Furthermore, considering narratology’s historical focus on long-form literary forms like the novel, Elinor Ochs ’s famous assertion that the “mundane conversational narratives of personal experience constitute the prototype of narrative activity rather than the flawed byproduct of more artful and planned narrative discourse” continues to highlight the opportunities for narrative theorists to attend to small stories, such as social media stories. We hope that the
@@
Compared to prior work that detects storytelling in online healthcare communities (Ganti et al., 2022, 2023), our model predicts an overall lower rate of storytelling, likely owing to the diversity of topics in our dataset. Our classification performance is lower than the performance reported in those works, despite using similar fine-tuning techniques, signalling the challenge of identifying stories across diverse topics and communities. While storytelling may be a sign of trust as a form of self-disclosure (Ma et al., 2019), we do not find a significant relationship between the the overall storytelling rate in a specific community and the same community’s toxicity (measured via the Perspective API555https://perspectiveapi.com/), size (number of members), or user activity (posts per member). Finally, we provide some additional experiments about overall post and comment patterns in Appendix A.4.
@@ 9 Conclusion
We formalize the task of story detection and introduce the task of story span detection by releasing
@@
StorySeeker, a framework that contains a dataset, codebook, and fine-tuned models.
@@ 11 Ethical Considerations
Our study was approved by the IRB at our institution.
@@ A.2 Additional Information About Dataset
gaming |
r/leagueoflegends, r/gaming, r/DotA2 |
hobbies |
r/poker, r/photography, r/MakeupAddiction |
tech |
r/technology, r/linux, r/AndroidQuestions |
fandom |
r/asoiaf, r/doctorwho, r/StarWars |
general |
r/AskReddit, r/pics, r/Showerthoughts |
informative |
r/explainlikeimfive, r/math, r/RealEstate |
news & politics |
r/PoliticalDiscussion, r/changemyview, r/Economics |
professional advice |
r/legaladvice, r/AskDocs, r/graphic_design |
professional sports |
r/nfl, r/hockey, r/LiverpoolFC |
relationships |
r/relationships, dating_advice, r/Parenting |
fitness |
r/running, r/bodybuilding, r/keto |
religion |
r/atheism, r/Christianity, r/DebateReligion |
stories |
r/tifu, r/TalesFromRetail, r/Dreams |
cities |
r/toronto, r/Seattle, r/LosAngeles |
countries |
r/canada, r/japan |
drugs |
r/Drugs, r/LSD, r/opiates |
mental health |
r/depression, r/ADHD, r/socialanxiety |
queer culture |
r/lgbt, r/ainbow, r/bisexual |
gender |
r/TwoXChromosomes, r/AskMen, r/AskWomen |
self help |
r/introvert, r/GetMotivated, r/INTP |
software dev |
r/programming, r/cscareerquestions, r/gamedev |
confessions |
r/offmychest, r/DoesAnybodyElse, r/confession |
finance |
r/personalfinance, r/Bitcoin, r/investing |
academic |
r/EngineeringStudents, r/college, r/GradSchool |
addiction |
stopdrinking, stopsmoking, cripplingalcoholism |
animals |
r/Pets, r/Dogtraining, r/cats |
| healthcare |
r/BabyBumps, r/diabetes, r/SkincareAddiction |
other |
banned subreddits, subreddits about children, NSFW and toxic subreddits, miscellaneous |
@@ A.3 Stories, Toics, and Winning Arguments in r/ChangeMyView
0 |
traffic accidents |
car, driving, cars, NUM, drive, traffic, stop, risk, dog, seat |
@@
Take a look at these statistics from Wikipedia: > According to the U.S. Natio… |
1 |
drug legalization |
@@
death, drugs, crime, society, health, alcohol, drug, illegal, consent, life |
The official death toll in the Mexican drug war states that sixty thousand peopl… |
2 |
economic systems |
@@
money, would, pay, people, government, work, tax, income, business, job |
## Section I: Why is Basic Income Increasingly Popular? "Basic income" is a poli… |
3 |
war & military |
war, us, country, states, government, military, gun, united, guns, countries |
@@
Let me be as clear as possible: I’m talking only about the Army, the main branch… |
4 |
time, wages, tips |
@@
NUM, time, year, sports, number, hour, every, team, three, football |
Here is an outline of my reasoning with MS paint diagrams: http://i.imgur.com/nY… |
5 |
ethical animal treatment |
animals, food, meat, eat, eating, animal, humans, vegetarian, healthy, diet |
@@
Given that livestock consume more water, feed, create more greenhouse gasses, an… |
6 |
mundane (food, hygiene) |
power, water, would, use, workers, gt, demand, paper, could, employers |
@@
Benn Wyatt made several good points about the virtues of calzones. For those tha… |
7 |
linguistics |
@@
culture, use, like, term, language, way, change, often, people, words |
Some cases in point: - The US/English pronunciation of the name Rothschild as "R… |
8 |
education |
@@
school, college, education, students, job, schools, work, student, high, learn |
Note: Although I do hold this opinion, it feels prejudiced to me so I am genuine… |
9 |
guilt & responsibility |
@@
people, person, someone, bad, think, kill, even, self, good, lives |
Batman’s strict no-killing policy has lead to the deaths of hundreds of people. … |
10 |
mental health & philosophy |
like, reddit, think, read, much, hate, seems, pretty, etc, mean |
@@
When discussing people in altered states, including those brought about through … |
11 |
rights & laws |
@@
right, rights, law, laws, free, gt, freedom, case, legal, argument |
I believe the modern Libertarian (as defined by people like Ron Paul) is hypocri… |
12 |
future population |
NUM, us, years, world, new, time, gt, better, population, means |
@@
*“Will robots inherit the earth? Yes, but they will be our children.” - Marvin M… |
13 |
work accountability |
people, would, work, go, get, life, time, could, much, day |
@@
I understand that child rearing is important, but how can I have equal respect f… |
14 |
films, games, art |
@@
game, games, art, play, video, movie, movies, show, character, characters |
I believe 3D is a cancer upon the film industry for the following reasons: * 3D … |
15 |
lifestyle choices |
even, many, go, end, ever, less, age, years, real, least |
@@
I have been drinking various craft beers for more than 20 years now. In the past… |
16 |
technology |
@@
use, ads, phone, buy, used, price, computer, store, internet, apple |
You probably get this a lot, but I was thinking about it in the shower today. Ye… |
17 |
abortion |
@@
child, children, parents, kids, abortion, life, mother, birth, pro, family |
Both the man and woman are equally responsible for an unplanned pregnancy. My re… |
18 |
music |
@@
music, great, even, time, like, sound, much, value, many, english |
Why do sound technicians ruin so many indoor gigs by turning the whole volume up… |
19 |
sexual violence |
@@
rape, victim, information, issues, victims, google, internet, see, reddit, sexual |
The argument from miracles is the argument that there have been miracles which a… |
20 |
social norms |
@@
like, people, get, want, really, know, see, re, think, would |
This is something that’s always bothered me, all the way back to when I was a ki… |
21 |
race |
@@
white, black, people, race, racism, racist, news, group, media, social |
I had a heated ideological debate last night with 2 sociologists and a feminist … |
22 |
global relations |
@@
society, many, state, world, believe, social, major, countries, science, philosophy |
It appears to me that when people talk about the rise of China as a global force… |
23 |
metaphysics |
@@
one, believe, think, based, evidence, matter, also, understanding, true, know |
First of all, the "you" you identify with is probably the summation of biologica… |
24 |
elections |
@@
vote, people, political, would, system, voting, party, politics, democracy, politicians |
People who have no interest in politics, and are not interested in learning abou… |
25 |
gender |
@@
women, men, sex, gender, male, gay, woman, man, female, sexual |
Gender, as defined by Wikipedia, is "the range of characteristics pertaining to,… |
26 |
words & symbols |
word, use, definition, gt, using, used, logic, said, first, considered |
@@
EDIT: Please refrain from making anymore comments about controlled demolitions, … |
27 |
meta-CMV |
@@
view, edit, people, think, would, change, changed, believe, point, post |
Lately (especially since the invasion from /r/adviceanimals), there have been a … |
28 |
space & evolution |
@@
would, one, could, life, likely, make, police, environment, might, human |
First off, let me clarify, I am not defending the actions of Ferguson Police Off… |
29 |
religion |
@@
god, religion, human, religious, believe, moral, good, would, beliefs, belief |
The typical Christian resolution to the problem of evil is to state that it is h… |

### v3

@@ 6.1 What are the signs of storytelling spans?
| Measure | Effect Size () | Direction | -value |
| expert-annotated events | 2.122*** | story | |
| past tense | 1.742*** | story | |
| realis events | 1.655*** | story | |
| 1st-person singular pronouns | 1.051*** | story | |
| 3rd-person singular pronouns | 0.459*** | story | |
| entity mentions | 0.346*** | story | |
| concreteness | 0.329*** | story | 0.001 |
| non-past tense | 1.296*** | non-story | |
| is comment (vs. post) | 0.612*** | non-story | |
| 2nd-person pronouns | 0.551*** | non-story | |
| sentence length | - | - | 0.345 |
| 1st-person plural pronouns | – | – | 0.345 |
Based on prior literature and our iterations of annotation and codebook refinement, we identify a set of features that we expect may be associated with storytelling in social media. These include entity and pronoun rates (Eisenberg and Finlayson, 2017; Piper and Bagga, 2022), events Hühn (2009); Gius and Vauth (2022); Sap et al. (2022); Sims et al. (2019), verb tense, concreteness (Piper and Bagga, 2022; Brysbaert et al., 2013), whether a text is a post or comment, and average sentence length. See A.9 for context on prior work related to these features, the precise definitions we adopted for this study, and a feature comparison test with another narrative detection dataset composed of mostly literary texts.
@@ 6.2 Where do people tell stories?
When interpreting these results, it is important to keep in mind that our comparisons are only across texts in the Webis-TLDR-17 dataset, i.e. coherent texts with summaries. Since many subreddits predominately include photos or other content, our storytelling rates should be interpreted as relative rankings rather than as absolute rates of storytelling. We provide additional results in Figure 5 in A.5 that are not restricted to texts with summaries, validating that general patterns hold when sampling over the full space of subreddit posts.
@@ 6.3 How distinctive are stories by community?
In Figure 2, we map communities across two axes: their story rate, predicted by our story detector, and the distinctiveness of their vocabulary. Categories of subreddits form interpretable clusters. We show 291 subreddits (those matching our filtering criteria, see §5), and Table 9 in A.5 shows the “corners” of this plot, i.e., the subreddits with the most and least storytelling and distinctiveness. For example, subreddits in the stories category, such as r/Glitch_in_the_Matrix, tend to have both high rates of storytelling and low distinctiveness — these subreddits elicit stories that do not use consistently distinctive language. Some categories contain a wide range, e.g., the technology category contains subreddits varying across half the distinctiveness range, from r/apple (not at all distinctive) to r/buildapc (very distinctive stories). In Figure 6 in A.5, we show the full ranking of subreddit categories, which can aid researchers in determining whether to build a custom detector for their target domain or rely on a mixture of finetuning data to align with the diverse storytelling in some categories.
@@ 7 Case Study: r/ChangeMyView
Figure 3 shows the post topics that elicit comments with more or less storytelling and whose storytelling comments are more or less likely to receive a winning “delta” point, indicating that the poster was convinced by the comment’s argument. Topics are trained on the posts via a latent Dirichlet allocation (LDA) (Blei et al., 2003) model (), and after examination of the highest probability words and documents, we assign each topic a descriptive name. The y-axis shows the difference between the mean post topic probabilities for storytelling and non-storytelling comments, while the x-axis shows the difference between the proportion of winning comments that include storytelling and the proportion of non-winning comments that include storytelling. More details about our methods and the topic model are given in A.3.
@@ 8 Discussion
The features that are most strongly associated with storytelling support prior work’s emphasis on storytelling’s basis in agent-centered, event-driven forms of communication that are grounded in concrete settings. These features are believed to support social coordination by fostering joint attention around virtual events (known as the “deictic theory” of storytelling; Piper and Bagga, 2022).
Nevertheless, considering narratology’s historical focus on long-form literary forms like the novel, Elinor Ochs ’s famous assertion that the “mundane conversational narratives of personal experience constitute the prototype of narrative activity rather than the flawed byproduct of more artful and planned narrative discourse” highlights the opportunities for narrative theorists to attend to “small stories” (Georgakopoulou, 2007), such as social media stories profiled here. We hope that the
@@
Compared to prior work that detects storytelling in online healthcare communities (Ganti et al., 2022, 2023), our model predicts an overall lower rate of storytelling, likely owing to the diversity of topics in our dataset. Our classification performance is lower than the performance reported in those works, despite using similar fine-tuning techniques, signalling the challenge of identifying stories across diverse topics and communities. While storytelling may be a sign of trust as a form of self-disclosure (Ma et al., 2019), we do not find a significant relationship between the overall storytelling rate in a specific community and the same community’s toxicity (measured via the Perspective API555https://perspectiveapi.com/), size (number of members), or user activity (posts per member). Finally, we provide some additional experiments about overall post and comment patterns in A.4.
@@ 9 Conclusion
We formalize the task of story detection and story span detection by releasing
@@
StorySeeker, a framework that contains a dataset, codebook, and fine-tuned models for online story detection.
@@ 10 Limitations
Our annotation procedure can include multiple contiguous spans of story spans for a given text. When a story is interrupted by non-story text, we highlight the story spans and do not highlight the interrupting non-story span. We observed many cases like this in the dataset, emphasizing the importance of our span annotations (rather than labeling the entire text with a single binary label). However, we do not capture whether non-contiguous spans are part of the same story. Future work could augment our annotations with such story span coreference information; this would be a valuable addition to our dataset.
@@ 11 Ethical Considerations
Our study was considered exempt by the IRB at the Allen Institute for AI.
@@ A.2 Additional Information About Dataset
| gaming | r/leagueoflegends, r/gaming, r/DotA2 |
| hobbies | r/poker, r/photography, r/MakeupAddiction |
| tech | r/technology, r/linux, r/AndroidQuestions |
| fandom | r/asoiaf, r/doctorwho, r/StarWars |
| general | r/AskReddit, r/pics, r/Showerthoughts |
| informative | r/explainlikeimfive, r/math, r/RealEstate |
| news & politics | r/PoliticalDiscussion, r/changemyview, r/Economics |
| professional advice | r/legaladvice, r/AskDocs, r/graphic_design |
| professional sports | r/nfl, r/hockey, r/LiverpoolFC |
| relationships | r/relationships, dating_advice, r/Parenting |
| fitness | r/running, r/bodybuilding, r/keto |
| religion | r/atheism, r/Christianity, r/DebateReligion |
| stories | r/tifu, r/TalesFromRetail, r/Dreams |
| cities | r/toronto, r/Seattle, r/LosAngeles |
| countries | r/canada, r/japan |
| drugs | r/Drugs, r/LSD, r/opiates |
| mental health | r/depression, r/ADHD, r/socialanxiety |
| queer culture | r/lgbt, r/ainbow, r/bisexual |
| gender | r/TwoXChromosomes, r/AskMen, r/AskWomen |
| self help | r/introvert, r/GetMotivated, r/INTP |
| software dev | r/programming, r/cscareerquestions, r/gamedev |
| confessions | r/offmychest, r/DoesAnybodyElse, r/confession |
| finance | r/personalfinance, r/Bitcoin, r/investing |
| academic | r/EngineeringStudents, r/college, r/GradSchool |
| addiction | stopdrinking, stopsmoking, cripplingalcoholism |
| animals | r/Pets, r/Dogtraining, r/cats |
| healthcare | r/BabyBumps, r/diabetes, r/SkincareAddiction |
| other | banned subreddits, subreddits about children, NSFW and toxic subreddits, miscellaneous |
@@ A.3 Stories, Toics, and Winning Arguments in r/ChangeMyView
| 0 | traffic accidents | car, driving, cars, NUM, drive, traffic, stop, risk, dog, seat | Take a look at these statistics from Wikipedia: > According to the U.S. Natio… |
@@
| 1 | drug legalization | death, drugs, crime, society, health, alcohol, drug, illegal, consent, life | The official death toll in the Mexican drug war states that sixty thousand peopl… |
@@
| 2 | economic systems | money, would, pay, people, government, work, tax, income, business, job | ## Section I: Why is Basic Income Increasingly Popular? "Basic income" is a poli… |
@@
| 3 | war & military | war, us, country, states, government, military, gun, united, guns, countries | Let me be as clear as possible: I’m talking only about the Army, the main branch… |
@@
| 4 | time, wages, tips | NUM, time, year, sports, number, hour, every, team, three, football | Here is an outline of my reasoning with MS paint diagrams: http://i.imgur.com/nY… |
@@
| 5 | ethical animal treatment | animals, food, meat, eat, eating, animal, humans, vegetarian, healthy, diet | Given that livestock consume more water, feed, create more greenhouse gasses, an… |
@@
| 6 | mundane (food, hygiene) | power, water, would, use, workers, gt, demand, paper, could, employers | Benn Wyatt made several good points about the virtues of calzones. For those tha… |
@@
| 7 | linguistics | culture, use, like, term, language, way, change, often, people, words | Some cases in point: - The US/English pronunciation of the name Rothschild as "R… |
@@
| 8 | education | school, college, education, students, job, schools, work, student, high, learn | Note: Although I do hold this opinion, it feels prejudiced to me so I am genuine… |
@@
| 9 | guilt & responsibility | people, person, someone, bad, think, kill, even, self, good, lives | Batman’s strict no-killing policy has lead to the deaths of hundreds of people. … |
@@
| 10 | mental health & philosophy | like, reddit, think, read, much, hate, seems, pretty, etc, mean | When discussing people in altered states, including those brought about through … |
@@
| 11 | rights & laws | right, rights, law, laws, free, gt, freedom, case, legal, argument | I believe the modern Libertarian (as defined by people like Ron Paul) is hypocri… |
@@
| 12 | future population | NUM, us, years, world, new, time, gt, better, population, means | *“Will robots inherit the earth? Yes, but they will be our children.” - Marvin M… |
@@
| 13 | work accountability | people, would, work, go, get, life, time, could, much, day | I understand that child rearing is important, but how can I have equal respect f… |
@@
| 14 | films, games, art | game, games, art, play, video, movie, movies, show, character, characters | I believe 3D is a cancer upon the film industry for the following reasons: * 3D … |
@@
| 15 | lifestyle choices | even, many, go, end, ever, less, age, years, real, least | I have been drinking various craft beers for more than 20 years now. In the past… |
@@
| 16 | technology | use, ads, phone, buy, used, price, computer, store, internet, apple | You probably get this a lot, but I was thinking about it in the shower today. Ye… |
@@
| 17 | abortion | child, children, parents, kids, abortion, life, mother, birth, pro, family | Both the man and woman are equally responsible for an unplanned pregnancy. My re… |
@@
| 18 | music | music, great, even, time, like, sound, much, value, many, english | Why do sound technicians ruin so many indoor gigs by turning the whole volume up… |
@@
| 19 | sexual violence | rape, victim, information, issues, victims, google, internet, see, reddit, sexual | The argument from miracles is the argument that there have been miracles which a… |
@@
| 20 | social norms | like, people, get, want, really, know, see, re, think, would | This is something that’s always bothered me, all the way back to when I was a ki… |
@@
| 21 | race | white, black, people, race, racism, racist, news, group, media, social | I had a heated ideological debate last night with 2 sociologists and a feminist … |
@@
| 22 | global relations | society, many, state, world, believe, social, major, countries, science, philosophy | It appears to me that when people talk about the rise of China as a global force… |
@@
| 23 | metaphysics | one, believe, think, based, evidence, matter, also, understanding, true, know | First of all, the "you" you identify with is probably the summation of biologica… |
@@
| 24 | elections | vote, people, political, would, system, voting, party, politics, democracy, politicians | People who have no interest in politics, and are not interested in learning abou… |
@@
| 25 | gender | women, men, sex, gender, male, gay, woman, man, female, sexual | Gender, as defined by Wikipedia, is "the range of characteristics pertaining to,… |
@@
| 26 | words & symbols | word, use, definition, gt, using, used, logic, said, first, considered | EDIT: Please refrain from making anymore comments about controlled demolitions, … |
@@
| 27 | meta-CMV | view, edit, people, think, would, change, changed, believe, point, post | Lately (especially since the invasion from /r/adviceanimals), there have been a … |
@@
| 28 | space & evolution | would, one, could, life, likely, make, police, environment, might, human | First off, let me clarify, I am not defending the actions of Ferguson Police Off… |
@@
| 29 | religion | god, religion, human, religious, believe, moral, good, would, beliefs, belief | The typical Christian resolution to the problem of evil is to state that it is h… |